OpenStudy (anonymous):

What is the slope of the line through the points (2/3,4/7) and (2/3,11/7) ?

OpenStudy (anonymous):

\[\huge\dfrac{ \dfrac{ 11 }{ 7 }-\dfrac{ 4 }{ 7 } }{ \dfrac{ 2 }{ 3 }-\dfrac{ 2 }{ 3 }}=\dfrac{ \dfrac{ 7 }{ 7 } }{ 0 }\] \[=\dfrac{1}{0}\] \[=undefined\]

OpenStudy (anonymous):

remember to always use the equation of a slope.